y - 1
------ = 1/5
x + 3
Solve for Y

Answer by wuwei96815(245)

y-1
----- = 1/5
x+3

Multiply both sides by x+3 gives:
y-1 = x+3/5
Multiply both sides by 5 gives:
5y-5 = x+3
Add 5 to each side:
5y = x+8
Divide both sides by 5:
y = x+8/5